Hello, could i get some help finding a solution to this?
Find all values satisfying the equation |x-9| = 3.
Enter your answers in increasing order.
x= ?
x= ?
thank you!

Answer by dkppathak(439)

Find all values satisfying the equation
Enter your answers in increasing order.|x-9| = 3.
x= ?
x= ?
solution
we know that
|x| =x when x is greater than 0 x>0
|x| =-x when x<0 using above property of modulus
|x-9| = 3
x-9=3 or -x+9=3
x=3+9 or -x=3-9
x=12 or -x=-6
x=12 or x=6